OpenStudy (anonymous):

What are the discontinuities of the function f(x) = the quantity x squared minus 16 over the quantity 4x plus 24 ?

OpenStudy (anonymous):

\[f(x) = (x-16)^2/(4x+24)\] so factor the numerator and denominator

OpenStudy (campbell_st):

well factor both the numerator and denominators

OpenStudy (anonymous):

\[\frac{ (x-4)(x+4) }{ (4)(x+6) }\]

OpenStudy (campbell_st):

\[\frac{x^2 - 16}{4x + 24}\] you will find a common factor.... but 1st solve 4x + 24 = 0 to find the discontinuity

OpenStudy (anonymous):

the denominator can't equal zero so set the denominator equal to 0

OpenStudy (anonymous):

4(x+6) = 0

OpenStudy (anonymous):

x = -6

OpenStudy (anonymous):

Wait so are you saying the discontinuity is -6?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

you're correct @nicjohns it is not 0... that's why you shouldn't post answers

OpenStudy (anonymous):

Thank you!!!